OpenStudy (anonymous):

find sec theta if tan theta =7/24 and cos theta < 0.

OpenStudy (campbell_st):

do you know which quadrant has tan as positive and cos negative..?

OpenStudy (anonymous):

quadrant 3

OpenStudy (campbell_st):

great so drawing a diagram for quadrant 3  can you now find c, the hypotenuse using pythagoras?

OpenStudy (anonymous):

he said do not draw a triangle to use the identities I think

OpenStudy (anonymous):

c=sqrt(a^2+b^2)

OpenStudy (campbell_st):

ok then so you need to know \[\sin^2(\theta) + \cos^1(\theta) = 1\] dividing every term by cos^2 gives \[\tan^2(\theta) + 1 = \sec^2(\theta)\] so substituting \[(\frac{7}{24})^2 + 1 = \sec^2(\theta)\] so evaluate this and then take the square root. then use the negative value.

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

thanks so much I just found where I wrote it all down in my notes !!!